# Universal Solvability for Robot Motion Planning on Graphs

### v1

@@ 2.2 Problem Definition
Yu and Rus [24] proposed a subroutine that solves the FRMP problem for a given instance consisting $(G(V,E),p,(S,T))$ in time $\mathcal{O}(|V|+|E|)$, based on the Theorem 20 in their paper.
We denote this algorithm as $\mathcal{A}$ in our work, and for the sake of completeness, we state Theorem 20 from their paper [24] again, as per the notations used in our paper.
@@ 3 Accumulating Robots to Specific Vertices
Please refer to Figure 4 for a visual representation of the execution of the algorithm. By construction, if $S^{\prime}$ is the output of Algorithm 1 when the configuration $S$ is part of the input, then $S^{\prime}(R)=V_{p}$. We now show that Algorithm 1 terminates in finite time.
Figure 4: Illustration of working of Algorithm 1, where $r_{i}$ denotes robot $i$.
@@
Algorithm 1 terminates in finite time.
@@
Let $d_{v}$ be the distance of a node $v$ to its nearest node in $V_{p}$ in the graph $\mathfrak{T}$. At each iteration, $d_{S(i)}$ can only decrease, for any $i\in R$. Moreover, there exists $i\in R$ such that $d_{S(i)}$ strictly decreases, because either some robot becomes included in $V_{p}$, or robots move along a path and are placed nearer to $V_{p}$. Therefore, the non-negative integer value $\sum\limits_{i\in R}d_{S(i)}$ strictly decreases at each iteration. Hence, Algorithm 1 must terminate in finite time.
We now argue that if the positions of the robots are permuted in $S$, the outputs of Algorithm 1 are permuted correspondingly.
@@
Consider a graph $G(V,E)$, a set of robots $R=[p]$ and an intrinsic ordering ${\sf Intr}$ of $V$. Let $S_{1}$ and $S_{2}$ be configurations such that they occupy the same set of vertices, i.e., $S_{1}(R)=S_{2}(R)$. Therefore, there exists a permutation $\pi:R\to R$ such that $S_{1}(i)=S_{2}(\pi(i))$ for all $i\in R$. Let $S^{\prime}_{1}$ be the output configuration when Algorithm 1 is run on $G$ and $S_{1}$; and let $S^{\prime}_{2}$ be the output configuration when Algorithm 1 is run on $G$ and $S_{2}$. Then for all $i\in R$, $S^{\prime}_{1}(i)=S^{\prime}_{2}(\pi(i))$
Therefore, if robot $r$ occupies vertex $v$ in $S$ i.e., $S(r)$, then the final position of $r$ in $S^{\prime}$ i.e., $S^{\prime}(r)$ is only dependent on $v$ and the set $S(R)$; where $S^{\prime}$ is the output of Algorithm 1 on $G$ and $S$. This motivates us to define the accumulation map.
@@
where $S^{\prime}$ is the output of Algorithm 1 on $G$ and $S$.
Thus, ${\mathcal{T}_{S(R)}}$ maps a set of occupied vertices $S(R)$ to $V_{p}$. From Lemma 3.3, we have ${\mathcal{T}_{S_{1}(R)}}$ and ${\mathcal{T}_{S_{2}(R)}}$ are identical whenever $S_{1}(R)=S_{2}(R)$, i.e., the configurations $S_{1}$ and $S_{2}$ occupy the same set of vertices. Therefore, for a $p$-sized subset $\alpha\in\binom{V}{p}$, this allows us to uniquely define ${\mathcal{T}_{\alpha}}:\alpha\to V_{p}$ to be equal to ${\mathcal{T}_{S(R)}}$ for any configuration $S$ with $S(R)=\alpha$. We can hence define the output $Y_{S}$ of Algorithm 1 when run on $G$ and $S$ as $Y_{S}={\mathcal{T}_{S(R)}}\circ S$, where $\circ$ is the composition function. Here, $S$ is a map from $R$ to $S(R)\subseteq V$, and ${\mathcal{T}_{S(R)}}$ is a map from $S(R)$ to $V_{p}$, hence $Y_{S}={\mathcal{T}_{S(R)}}\circ S$ is a map from $R$ to $V_{p}$.
@@
Consider a graph $G(V,E)$, a set of robots $R=[p]$ and an intrinsic ordering ${\sf Intr}$ of $V$. For any configuration $S$, we say the configuration $Y_{S}={\mathcal{T}_{S(R)}}\circ S$, the output of Algorithm 1 on $S$, is the accumulation of $S$.
Since Algorithm 1 performs valid moves, we have the following observation.
@@
If $S$ is a configuration such that $S(R)=V_{p}$, the Algorithm 1 stops without entering the loop at Line 4. Therefore, we get $Y_{S}=S$.
@@
Recall that $(S,S^{\prime})$ is a valid move if and only if $(S^{\prime},S)$ is a valid move. Moreover, if $S$ and $T$ are configurations, then $S\sim Y_{S}$ and $T\sim Y_{T}$ (Section 3). This implies the following result.
@@ 4 Randomized Algorithm for USolR
In this section, we design a polynomial-time randomized algorithm for the USolR problem based on the sizes of the equivalence classes of the relation $\tilde{R}$ (Definition 2.2). The notions of the reachability relation and equivalence classes have resemblance to groups and subgroups. These similarities, as well as the group-theoretic techniques used in the works of Yu and Rus [24] motivated us to deeply examine the structure of these equivalence classes of the reachability relation $\tilde{R}$, and we provide a characterisation of the equivalence classes of $\tilde{R}$.
@@ 4.1 Equivalence Classes of the Reachability Relation
Let $(G(V,E),p)$ be the instance of USolR, and let $V_{p}=\{v_{1},v_{2},\ldots,v_{p}\}$ be as defined in Section 3. Consider the identity configuration $S_{I}$ with $S_{I}(i)=v_{i}$ for all $i\in[p]$. For configuration $S_{I}$, we have $S_{I}(R)=V_{p}$, so $Y_{S_{I}}=S_{I}$ (Section 3).
For a graph $G(V,E)$ to be universally solvable for $p$ robots, it must hold true that for any pair of configurations $S,T\in\mathcal{S}_{G}$, $T$ is reachable from $S$, i.e., $(S,T)\in\tilde{R}$. Thus, for any configuration $S\in\mathcal{S}_{G}$, $S_{I}\sim S$ must also hold true. Further, from the discussion in Section 2, we have that $\tilde{R}$ is an equivalence relation, and since all configurations in $\mathcal{S}_{G}$ belong to the same equivalence class as $S_{I}$, there is essentially only one equivalence class comprising of all configurations $S\in\mathcal{S}_{G}$. Let us denote by $\mathcal{E}_{S}$, the set of configurations reachable from the configuration $S$. Then, for a YES-instance, we have $\mathcal{E}_{S_{I}}=\mathcal{S}_{G}$.
@@
Consider a graph $G(V,E)$, a set of robots $R=[p]$ and an intrinsic ordering ${\sf Intr}$ of $V$. Let $\tilde{R}$ be the equivalence relation on configurations as per Definition 2.2. Then the size of the equivalence class $\mathcal{E}_{S_{I}}$ equals the size of the equivalence class $\mathcal{E}_{S_{j}^{*}}$ for all $j\in\{1,2,\ldots,k\}$.
@@
We have $S_{I}\not\sim S_{j}^{*}$. Let $Y_{S_{j}^{*}}$ denote the configuration obtained by running Algorithm 1 on $S_{1}^{*}$ (Definition 3.5). By Section 3, $S_{j}^{*}\sim Y_{S_{j}^{*}}$. Therefore, $S_{I}\not\sim Y_{S_{j}^{*}}$ and $Y_{S_{j}^{*}}\in\mathcal{E}_{S_{j}^{*}}$. Since $Y_{S_{j}^{*}}$ is a bijection from $R$ to $V_{p}$, there exists a permutation $\pi^{*}$ such that $Y_{S_{j}^{*}}=S_{I}\circ\pi^{*}$ (i.e., $Y_{S_{j}^{*}}(i)=S_{I}(\pi^{*}(i))=v_{\pi^{*}(i)}$ for all $i\in[p]$). Next, we define a function $f_{\pi^{*}}:\mathcal{E}_{S_{I}}\to\mathcal{E}_{S_{j}^{*}}$ as $f_{\pi^{*}}(S)=S\circ\pi^{*}$, for all $S\in\mathcal{E}_{S_{I}}$. Note that the function $f_{\pi^{*}}$ is defined according to the permutation $\pi^{*}$ that relates $Y_{S_{j}^{*}}$ to $S_{I}$. We first establish that $f_{\pi^{*}}$ maps all configurations in $\mathcal{E}_{S_{I}}$ only to configurations in $\mathcal{E}_{S_{j}^{*}}$. With this, for $S_{I}$, we know that $f_{\pi^{*}}(S_{I})=Y_{S_{j}^{*}}\in\mathcal{E}_{S_{j}^{*}}$. For any other configuration $\tilde{S}\in\mathcal{E}_{S_{I}}$, there exists a sequence of configurations $S_{1}(=S_{I}),S_{2},...,S_{h}(=\tilde{S})$, where $(S_{t},S_{t+1})$ is a valid move for every $t\in[h-1]$. Now, consider the sequence of configurations $f_{\pi^{*}}(S_{1})=S_{1}\circ\pi^{*}=Y_{S_{j}^{*}},f_{\pi^{*}}(S_{2})=S_{2}\circ\pi^{*},...,f_{\pi^{*}}(S_{h})=\tilde{S}\circ\pi^{*}$. In this sequence, $(f_{\pi^{*}}(S_{t}),f_{\pi^{*}}(S_{t+1}))$ is a valid move for every $t\in[h-1]$ (Section 2). This corresponds to a sequence of valid moves from $Y_{S_{j}^{*}}$ to $f_{\pi^{*}}(\tilde{S})$. Since the argument is independent of the configuration $\tilde{S}\in\mathcal{E}_{S_{I}}$, we have that $f_{\pi^{*}}(\tilde{S})\in\mathcal{E}_{S_{j}^{*}}$, for all $\tilde{S}\in\mathcal{E}_{S_{I}}$. Thus, we have established that $f_{\pi^{*}}$ maps $\tilde{S}\in\mathcal{E}_{S_{I}}$ to configurations only in $\mathcal{E}_{S_{j}^{*}}$.
@@
From Lemma 4.1, we have $|\mathcal{E}_{S_{I}}|=|\mathcal{E}_{S_{i}^{*}}|$ for all $i\in[k]$, and also $|\mathcal{S}_{G}|=\bigcup\limits_{i=0}^{k}|\mathcal{E}_{S_{i}^{*}}|$. Further, since the equivalence classes induce a partition over $\mathcal{S}_{G}$, we have $|\mathcal{S}_{G}|=\sum\limits_{i=0}^{k}|\mathcal{E}_{S_{i}^{*}}|$. Therefore, the size of each equivalence class is $|\mathcal{E}_{S_{i}^{*}}|=|\mathcal{S}_{G}|/(k+1)$ for all $i\in\{0,1,..,k\}$. From this we can conclude that if for a given NO-instance $(G(V,E),p)$, the relation $\tilde{R}$ is partitioned into $k+1$ equivalence classes $\mathcal{E}_{S_{0}^{*}},\mathcal{E}_{S_{1}^{*}},..,\mathcal{E}_{S_{k}^{*}}$ of equal sizes, then the configuration $S_{I}$ is not reachable to $(k\cdot|\mathcal{S}_{G}|)/(k+1)$ many configurations in $\mathcal{S}_{G}$. Since $k\geq 1$ for a NO-instance, $S_{I}$ is not reachable to $(k\cdot|\mathcal{S}_{G}|)/(k+1)\geq|\mathcal{S}_{G}|/2$ many configurations in $\mathcal{S}_{G}$.
@@ 4.2 Designing a randomized algorithm
We now propose a randomized algorithm for USolR based on the properties of equivalence classes discussed in Section 4.1. Given a USolR instance $(G(V,E),p)$, the randomized algorithm correctly outputs YES for a YES-instance (universally solvable instance) with probability $1$ and outputs YES for a NO-instance (not universally solvable instance) with probability at most $1/2$. We state the following theorem.
@@
Given an instance $(G(V,E),p)$ for USolR, we propose a randomized algorithm, denoted by Algorithm 2. Algorithm 2 first finds the connected components of the given graph $G$ using a Depth-First Search algorithm. If there are $2$ or more connected components in $G$, then it outputs NO and terminates, as a disconnected graph is never universally solvable. Next, Algorithm 2 samples a configuration $S_{R}$ uniformly at random from the set of all configurations $\mathcal{S}_{G}=\{S:R\to V\mid S\text{ is a configuration}\}$. Then, it runs Algorithm $\mathcal{A}$ (Proposition 2.4) as a subroutine, with $(G(V,E),p,(S_{I},S_{R}))$ as input to check if $S_{R}$ is reachable from $S_{I}$. If Algorithm $\mathcal{A}$ outputs YES for reachability of $S_{R}$ from $S_{I}$, then Algorithm 2 outputs that $G$ is universally solvable for $p$ robots (i.e., $(G(V,E),p)$ is a YES-instance), otherwise it outputs that $G$ is not universally solvable for $p$ robots (i.e., $(G(V,E),p)$ is a NO-instance). The pseudocode of Algorithm 2 can be found in the Appendix.
For a YES-instance $(G(V,E),p)$, all configurations $S\in\mathcal{S}_{G}$ are reachable from $S_{I}$. Therefore, when Algorithm 2 runs on a YES-instance, for any uniformly sampled configuration $S_{R}$, the feasibility subroutine Algorithm $\mathcal{A}$ will always return that $S_{R}$ is reachable from $S_{I}$, and hence Algorithm 2 would return YES with probability $1$.
For a NO-instance, the number of distinct configurations that are reachable from $S_{I}$ is at most $(|\mathcal{S}_{G}|/2)$ (Lemma 4.3). Algorithm 2 outputs YES for a NO-instance if Algorithm $\mathcal{A}$ outputs that $S_{R}$ (the uniformly sampled configuration) is reachable from $S_{I}$. Therefore, the probability that a uniformly sampled configuration $S_{R}$ is reachable from $S_{I}$ is equal to the ratio of the number of configurations in $\mathcal{S}_{G}$ reachable from $S_{I}$ to the total number of configurations in $\mathcal{S}_{G}$.
@@
We now analyse the time complexity of Algorithm 2. Sampling a random configuration $S_{R}$ from the set of all possible configurations can be done in linear time, $\mathcal{O}(|V|+|E|)$. Algorithm $\mathcal{A}$ runs in linear time, $\mathcal{O}(|V|+|E|)$ (Proposition 6.1). Moreover, all other steps (taking input, running DFS, etc.) are at most linear time, $\mathcal{O}(|V|+|E|)$ operations. Hence, Algorithm 2 runs in time $\mathcal{O}(|V|+|E|)$. This completes the proof of the theorem.

### v2

@@ 2.2 Problem Definition
Yu and Rus [yu2014pebblemotiongraphsrotations] proposed a subroutine that solves the FRMP problem for a given instance consisting $(G(V,E),p,(S,T))$ in time $\mathcal{O}(|V|+|E|)$, based on the Theorem 20 in their paper.
We denote this algorithm as $\mathcal{A}$ in our work, and for the sake of completeness, we state Theorem 20 from their paper [yu2014pebblemotiongraphsrotations] again, as per the notations used in our paper.
@@ 3 Accumulating Robots to Specific Vertices
Please refer to Figure˜4 for a visual representation of the execution of the algorithm. By construction, if $S^{\prime}$ is the output of Algorithm˜1 when the configuration $S$ is part of the input, then $S^{\prime}(R)=V_{p}$. We now show that Algorithm˜1 terminates in finite time.
Figure 4: Illustration of working of Algorithm˜1, where $r_{i}$ denotes robot $i$.
@@
Algorithm˜1 terminates in finite time.
@@
Let $d_{v}$ be the distance of a node $v$ to its nearest node in $V_{p}$ in the graph $\mathfrak{T}$. At each iteration, $d_{S(i)}$ can only decrease, for any $i\in R$. Moreover, there exists $i\in R$ such that $d_{S(i)}$ strictly decreases, because either some robot becomes included in $V_{p}$, or robots move along a path and are placed nearer to $V_{p}$. Therefore, the non-negative integer value $\sum\limits_{i\in R}d_{S(i)}$ strictly decreases at each iteration. Hence, Algorithm˜1 must terminate in finite time.
We now argue that if the positions of the robots are permuted in $S$, the outputs of Algorithm˜1 are permuted correspondingly.
@@
Consider a graph $G(V,E)$, a set of robots $R=[p]$ and an intrinsic ordering $\sf Intr$} of $V$. Let $S_{1}$ and $S_{2}$ be configurations such that they occupy the same set of vertices, i.e., $S_{1}(R)=S_{2}(R)$. Therefore, there exists a permutation $\pi:R\to R$ such that $S_{1}(i)=S_{2}(\pi(i))$ for all $i\in R$. Let $S^{\prime}_{1}$ be the output configuration when Algorithm 1 is run on $G$ and $S_{1}$; and let $S^{\prime}_{2}$ be the output configuration when Algorithm˜1 is run on $G$ and $S_{2}$. Then for all $i\in R$, $S^{\prime}_{1}(i)=S^{\prime}_{2}(\pi(i))$
Therefore, if robot $r$ occupies vertex $v$ in $S$ i.e., $S(r)$, then the final position of $r$ in $S^{\prime}$ i.e., $S^{\prime}(r)$ is only dependent on $v$ and the set $S(R)$; where $S^{\prime}$ is the output of Algorithm˜1 on $G$ and $S$. This motivates us to define the accumulation map.
@@
where $S^{\prime}$ is the output of Algorithm˜1 on $G$ and $S$.
Thus, $\mathcal{T}_{S(R)}$} maps a set of occupied vertices $S(R)$ to $V_{p}$. From Lemma˜3.3, we have $\mathcal{T}_{S_{1}(R)}$} and $\mathcal{T}_{S_{2}(R)}$} are identical whenever $S_{1}(R)=S_{2}(R)$, i.e., the configurations $S_{1}$ and $S_{2}$ occupy the same set of vertices. Therefore, for a $p$-sized subset $\alpha\in\binom{V}{p}$, this allows us to uniquely define ${\mathcal{T}_{\alpha}}:\alpha\to V_{p}$ to be equal to $\mathcal{T}_{S(R)}$} for any configuration $S$ with $S(R)=\alpha$. We can hence define the output $Y_{S}$ of Algorithm˜1 when run on $G$ and $S$ as $Y_{S}={\mathcal{T}_{S(R)}}\circ S$, where $\circ$ is the composition function. Here, $S$ is a map from $R$ to $S(R)\subseteq V$, and $\mathcal{T}_{S(R)}$} is a map from $S(R)$ to $V_{p}$, hence $Y_{S}={\mathcal{T}_{S(R)}}\circ S$ is a map from $R$ to $V_{p}$.
@@
Consider a graph $G(V,E)$, a set of robots $R=[p]$ and an intrinsic ordering $\sf Intr$} of $V$. For any configuration $S$, we say the configuration $Y_{S}={\mathcal{T}_{S(R)}}\circ S$, the output of Algorithm˜1 on $S$, is the accumulation of $S$.
Since Algorithm˜1 performs valid moves, we have the following observation.
@@
If $S$ is a configuration such that $S(R)=V_{p}$, the Algorithm˜1 stops without entering the loop at Line 4. Therefore, we get $Y_{S}=S$.
@@
Recall that $(S,S^{\prime})$ is a valid move if and only if $(S^{\prime},S)$ is a valid move. Moreover, if $S$ and $T$ are configurations, then $S\sim Y_{S}$ and $T\sim Y_{T}$ (Section˜3). This implies the following result.
@@ 4 Randomized Algorithm for USolR
In this section, we design a polynomial-time randomized algorithm for the USolR problem based on the sizes of the equivalence classes of the relation $\tilde{R}$ (Definition˜2.2). The notions of the reachability relation and equivalence classes have resemblance to groups and subgroups. These similarities, as well as the group-theoretic techniques used in the works of Yu and Rus [yu2014pebblemotiongraphsrotations] motivated us to deeply examine the structure of these equivalence classes of the reachability relation $\tilde{R}$, and we provide a characterisation of the equivalence classes of $\tilde{R}$.
@@ 4.1 Equivalence Classes of the Reachability Relation
Let $(G(V,E),p)$ be the instance of USolR, and let $V_{p}=\{v_{1},v_{2},\ldots,v_{p}\}$ be as defined in Section 3. Consider the identity configuration $S_{I}$ with $S_{I}(i)=v_{i}$ for all $i\in[p]$. For configuration $S_{I}$, we have $S_{I}(R)=V_{p}$, so $Y_{S_{I}}=S_{I}$ (Section˜3).
For a graph $G(V,E)$ to be universally solvable for $p$ robots, it must hold true that for any pair of configurations $S,T\in\mathcal{S}_{G}$, $T$ is reachable from $S$, i.e., $(S,T)\in\tilde{R}$. Thus, for any configuration $S\in\mathcal{S}_{G}$, $S_{I}\sim S$ must also hold true. Further, from the discussion in Section˜2, we have that $\tilde{R}$ is an equivalence relation, and since all configurations in $\mathcal{S}_{G}$ belong to the same equivalence class as $S_{I}$, there is essentially only one equivalence class comprising of all configurations $S\in\mathcal{S}_{G}$. Let us denote by $\mathcal{E}_{S}$, the set of configurations reachable from the configuration $S$. Then, for a YES-instance, we have $\mathcal{E}_{S_{I}}=\mathcal{S}_{G}$.
@@
Consider a graph $G(V,E)$, a set of robots $R=[p]$ and an intrinsic ordering $\sf Intr$} of $V$. Let $\tilde{R}$ be the equivalence relation on configurations as per Definition˜2.2. Then the size of the equivalence class $\mathcal{E}_{S_{I}}$ equals the size of the equivalence class $\mathcal{E}_{S_{j}^{*}}$ for all $j\in\{1,2,\ldots,k\}$.
@@
[…]
@@
From Lemma˜4.1, we have $|\mathcal{E}_{S_{I}}|=|\mathcal{E}_{S_{i}^{*}}|$ for all $i\in[k]$, and also $|\mathcal{S}_{G}|=\bigcup\limits_{i=0}^{k}|\mathcal{E}_{S_{i}^{*}}|$. Further, since the equivalence classes induce a partition over $\mathcal{S}_{G}$, we have $|\mathcal{S}_{G}|=\sum\limits_{i=0}^{k}|\mathcal{E}_{S_{i}^{*}}|$. Therefore, the size of each equivalence class is $|\mathcal{E}_{S_{i}^{*}}|=|\mathcal{S}_{G}|/(k+1)$ for all $i\in\{0,1,..,k\}$. From this we can conclude that if for a given NO-instance $(G(V,E),p)$, the relation $\tilde{R}$ is partitioned into $k+1$ equivalence classes $\mathcal{E}_{S_{0}^{*}},\mathcal{E}_{S_{1}^{*}},..,\mathcal{E}_{S_{k}^{*}}$ of equal sizes, then the configuration $S_{I}$ is not reachable to $(k\cdot|\mathcal{S}_{G}|)/(k+1)$ many configurations in $\mathcal{S}_{G}$. Since $k\geq 1$ for a NO-instance, $S_{I}$ is not reachable to $(k\cdot|\mathcal{S}_{G}|)/(k+1)\geq|\mathcal{S}_{G}|/2$ many configurations in $\mathcal{S}_{G}$.
@@ 4.2 Designing a randomized algorithm
We now propose a randomized algorithm for USolR based on the properties of equivalence classes discussed in Section˜4.1. Given a USolR instance $(G(V,E),p)$, the randomized algorithm correctly outputs YES for a YES-instance (universally solvable instance) with probability $1$ and outputs YES for a NO-instance (not universally solvable instance) with probability at most $1/2$. We state the following theorem.
@@
Given an instance $(G(V,E),p)$ for USolR, we propose a randomized algorithm, denoted by Algorithm˜2. Algorithm˜2 first finds the connected components of the given graph $G$ using a Depth-First Search algorithm. If there are $2$ or more connected components in $G$, then it outputs NO and terminates, as a disconnected graph is never universally solvable. Next, Algorithm˜2 samples a configuration $S_{R}$ uniformly at random from the set of all configurations $\mathcal{S}_{G}=\{S:R\to V\mid S\text{ is a configuration}\}$. Then, it runs Algorithm $\mathcal{A}$ (Proposition˜2.4) as a subroutine, with $(G(V,E),p,(S_{I},S_{R}))$ as input to check if $S_{R}$ is reachable from $S_{I}$. If Algorithm $\mathcal{A}$ outputs YES for reachability of $S_{R}$ from $S_{I}$, then Algorithm˜2 outputs that $G$ is universally solvable for $p$ robots (i.e., $(G(V,E),p)$ is a YES-instance), otherwise it outputs that $G$ is not universally solvable for $p$ robots (i.e., $(G(V,E),p)$ is a NO-instance). The pseudocode of Algorithm˜2 can be found in the Appendix.
For a YES-instance $(G(V,E),p)$, all configurations $S\in\mathcal{S}_{G}$ are reachable from $S_{I}$. Therefore, when Algorithm˜2 runs on a YES-instance, for any uniformly sampled configuration $S_{R}$, the feasibility subroutine Algorithm $\mathcal{A}$ will always return that $S_{R}$ is reachable from $S_{I}$, and hence Algorithm˜2 would return YES with probability $1$.
For a NO-instance, the number of distinct configurations that are reachable from $S_{I}$ is at most $(|\mathcal{S}_{G}|/2)$ (Lemma˜4.3). Algorithm˜2 outputs YES for a NO-instance if Algorithm $\mathcal{A}$ outputs that $S_{R}$ (the uniformly sampled configuration) is reachable from $S_{I}$. Therefore, the probability that a uniformly sampled configuration $S_{R}$ is reachable from $S_{I}$ is equal to the ratio of the number of configurations in $\mathcal{S}_{G}$ reachable from $S_{I}$ to the total number of configurations in $\mathcal{S}_{G}$.
@@
We now analyse the time complexity of Algorithm˜2. Sampling a random configuration $S_{R}$ from the set of all possible configurations can be done in linear time, $\mathcal{O}(|V|+|E|)$. Algorithm $\mathcal{A}$ runs in linear time, $\mathcal{O}(|V|+|E|)$ (Proposition˜6.1). Moreover, all other steps (taking input, running DFS, etc.) are at most linear time, $\mathcal{O}(|V|+|E|)$ operations. Hence, Algorithm˜2 runs in time $\mathcal{O}(|V|+|E|)$. This completes the proof of the theorem.